### v1

# Latent Self-Consistency for Reliable Majority-Set Selection in Short- and Long-Answer Reasoning

Abstract
Probabilistic decoding in Large Language Models (LLMs) often yields inconsistent outputs, particularly on complex or long-form questions. Self-Consistency (SC) mitigates this for short-form QA by majority voting over exact strings, whereas Universal Self-Consistency (USC) and Weighted Unigram Consistency Score (WUCS) extend to long-form responses but lose accuracy on short-form benchmarks.
We introduce Latent Self-Consistency (LSC), which selects the most semantically consistent response using learnable token embeddings. A lightweight forward generation of summary tokens increases inference time by less than 1% and requires no changes to the model architecture.
Across 6 short-form and 5 long-form reasoning benchmarks (e.g., MATH, MMLU, TruthfulQA), LSC surpasses SC, USC and WUCS on all short-form and long-form ones on average, while maintaining negligible computational overhead. These results position LSC as a practical consistency-selection method that works reliably across answer formats. Additionally, LSC provides well-calibrated confidence estimates, maintaining low Expected Calibration Error across both answer formats.
1 Introduction
Large Language Models (LLMs) have achieved remarkable success across diverse natural language processing tasks, from mathematical reasoning to code generation Brown et al. (2020); Chowdhery et al. (2023); OpenAI (2023). However, as inherently probabilistic generators, they frequently yield divergent outputs on repeated runs, undermining consistency and reliability in downstream applications. To mitigate this variability, Self-Consistency (SC) Wang et al. (2022) generates multiple candidate responses through independent sampling and applies majority voting to select the most frequent answer. This mechanism has proven highly effective, achieving substantial accuracy improvements in mathematical reasoning and multiple-choice benchmarks by filtering out anomalous reasoning paths.
However, SC’s reliance on exact string matching severely limits its applicability. While effective for "short-answer" tasks such as numerical problems or multiple-choice questions where alternative surface forms are rare, SC fails in "long-answer" scenarios including code generation, summarization, or detailed explanations, where semantically equivalent responses often differ lexically and structurally. To address this limitation, recent consistency-based extensions have emerged. Universal Self-Consistency (USC) Chen et al. (2024) employs the LLM itself as a semantic judge by concatenating all $N$ candidates into a single prompt for evaluation; however, this approach incurs substantial computational overhead (approximately 10% additional inference time) and significant memory usage (around 15% additional memory), while introducing potential judge-model biases. Weighted Unigram Consistency Score (WUCS) Jain et al. (2024) offers a lightweight alternative using term-frequency weighted Jaccard similarity with probabilistic weighting, but sacrifices accuracy on short-answer tasks and fails to capture semantic coherence in complex reasoning chains.
In this work, we introduce Latent Self-Consistency (LSC), a novel framework that achieves the best of both worlds: matching SC’s high accuracy on short-answer tasks while extending robustly to long-answer scenarios, all with minimal computational overhead (less than 1% additional inference time and negligible memory usage). LSC appends a small set of trainable tokens after the EOS token of each generated response and extracts their final-layer representations, enabling efficient KV cache reuse, to obtain compact embeddings that capture the semantic essence of each response. By computing pairwise cosine similarities in this learned embedding space, LSC identifies the most consistent response with negligible computational overhead and no complex text processing.
Our contributions are fourfold:
  • Universal majority selection across answer formats. We present the first consistency-based selection method that achieves competitive or superior performance on both short-answer tasks (matching SC) and long-answer tasks (surpassing USC and WUCS), without specific mode change per answer format.
  • Efficient semantic consistency via learned suffix embeddings. By encoding response semantics into compact learned embeddings, LSC achieves USC-level semantic understanding with significantly reduced latency and memory overhead, making it practical for real-time deployment.
  • Reliable and calibrated confidence estimates across tasks. LSC enables trustworthy uncertainty quantification, achieving low Expected Calibration Error (ECE) scores for both short-answer and long-answer benchmarks.
  • Dynamic boundary detection for robust majority set identification. We introduce a dynamic Top-$K$ boundary detection algorithm that automatically filters out noisy outliers, further enhancing selection accuracy by focusing on the most semantically coherent subset of responses.
Figure 1: Overview of the Latent Self-Consistency (LSC) inference framework. Step 1: Multiple candidate responses are generated. Step 2: A sequence of summary tokens is appended to each response, and a lightweight pass through the LLM yields the last hidden states of these tokens, which are averaged to produce a single representation per response. Step 3: Pairwise cosine similarities between these response representations are computed, and the candidate with the highest mean cosine similarity to all others is selected as the most consistent output.
2 Related Works
2.1 Consistency-Based Selection in LLMs
Self-Consistency (SC) Wang et al. (2023) generates $N$ independent samples and selects the most frequent answer via exact string matching, yielding strong gains on "short answer form" tasks but failing when semantically equivalent outputs vary lexically. Universal Self-Consistency (USC) Chen et al. (2024) uses the LLM itself as a referee by concatenating all $N$ candidates into a single prompt and performing one additional forward pass. However, this approach requires constructing entirely new prompts, preventing KV cache reuse and introducing substantial latency and potential judge-model bias. Unigram Consistency Score (UCS) and its weighted variant WUCS Jain et al. (2024) compute probability-weighted overlap measures for text similarity assessment, which while efficient, struggle to capture deep chain-of-thought coherence. Fine-grained Self-Consistency (FSC) Wang et al. (2024) integrates segment-level consensus by re-generating and synthesizing responses from scratch, incurring extra latency due to these additional generative passes. In contrast, our Latent Self-Consistency (LSC) extracts small, trainable token embeddings by simply appending learnable tokens after the EOS token of each generated response, enabling KV cache reuse for highly efficient forward passes. This approach unifies SC’s accuracy on short-answer tasks with robust, low-latency selection for long-answer tasks.
2.2 Prompt Tuning and Learnable Token Embeddings
Existing prompt tuning approaches Lester et al. (2021); Li and Liang (2021) typically append learnable tokens to user prompts for domain adaptation without full model fine-tuning. These methods operate in a pre-generation manner, modifying the input before the model generates responses.
Our work introduces the first approach to learn semantic summary representations of generated responses through learnable token embeddings. Our LSC method operates in a post-generation manner, appending learnable tokens after the assistant’s response generation is complete to capture the semantic essence of each response. We refer to these as "summary tokens" since they are trained to encode compact semantic summaries of the full responses. Crucially, since we only update the summary token embeddings while keeping the base LLM frozen, our method preserves the model’s original natural language capabilities without any degradation. This design enables efficient semantic consistency evaluation through lightweight forward passes while maintaining complete separation between consistency measurement and the model’s core competencies.
3 Method
Overview. Figure 1 illustrates the Latent Self-Consistency (LSC) inference pipeline in three sequential phases. Step 1, candidate generation: we prompt the LLM to produce $N$ responses. Step 2, post-generation encoding: we append $K$ learnable summary tokens after EoS token to each response and perform a lightweight forward pass through the LLM to yield final-layer representations of these tokens; these representations are averaged and normalized to form one compact representation per response. Step 3, majority answer detection: pairwise cosine similarities among representations are computed to identify and select the most representative response that aligns with the semantic majority.
We now describe the implementation of this pipeline through three key components: training procedure for learning semantic representations, efficient inference procedure, and majority answer selection.
3.1 Training Phase
3.1.1 Dataset Construction
We assemble a balanced mix of short-answer and long-answer samples using LLM-based pseudo-labeling. For each question, an LLM generates an answer from which we extract the answer string as the label, identifying which responses converge on the same conclusion. Short-answer data are drawn from GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021b), and TriviaQA Joshi et al. (2017). Long-answer samples are derived from MMLU Hendrycks et al. (2021a) by prompting chain-of-thought rationales and removing the final option letter to yield free-form explanations with known pseudo-labels. We maintain a 50:50 ratio of short to long formats. Detailed dataset construction procedures, statistics, and examples are provided in Appendix A, with prompt templates in Appendix B.
3.1.2 Learning representations via Supervised Contrastive Learning
To bring semantically consistent responses closer together, we use supervised contrastive learning: embeddings of responses sharing the same answer label are pulled together, while those with different labels are pushed apart. Concretely, we append $K$ learnable special tokens after each response and fine-tune only their embeddings (e.g. $K=6$, $d=4096$ for 8B LLM, updating $3\times 10^{-6}$ of parameters). Let $\{h_{i,1},\dots,h_{i,K}\}$ be the final-layer representations of those tokens for response $i$; we compute
$\bar{h}_{i}=\frac{1}{K}\sum_{m=1}^{K}h_{i,m},\quad z_{i}=\frac{\bar{h}_{i}}{\|\bar{h}_{i}\|}.$
With $y_{i}$ the answer label, the supervised contrastive loss He et al. (2020); Khosla et al. (2020) over a batch of size $N$ is
$\mathcal{L}_{\mathrm{SCL}}=-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{I}_{i}^{+}|}\sum_{j\in\mathcal{I}_{i}^{+}}\log\frac{\exp(z_{i}\!\cdot\!z_{j}/\tau)}{\sum_{k\neq i}\exp(z_{i}\!\cdot\!z_{k}/\tau)},$
where $\mathcal{I}_{i}^{+}=\{j\neq i:y_{j}=y_{i}\}$ and $\tau=0.07$.
3.2 Inference Phase
At inference, each candidate is appended with the same $K$ tokens after the EoS token and encoded in one lightweight forward pass to produce representations $\{z_{i}\}$. Unlike USC which concatenates all responses into new prompts requiring full re-computation, LSC appends tokens directly after the EoS token of each generated response, enabling reuse of the existing key-value (KV) cache from the generation phase. This means the forward pass only needs to compute representations for the $K$ additional tokens rather than processing entire sequences from scratch, resulting in minimal computational overhead (less than 1% latency increase and negligible memory usage). We then compute the cosine similarity matrix
$S_{ij}=\frac{z_{i}\cdot z_{j}}{\|z_{i}\|\|z_{j}\|}.$
3.3 Majority Answer Selection Phase
Our goal is to select the majority answer by identifying the response that has the highest average similarity to all other responses, effectively choosing the response that best represents the semantic consensus.
3.3.1 Exponentially-Weighted Mean Selection
We score each response by
$w_{i}=\frac{1}{N-1}\sum_{j\neq i}\exp\bigl{(}S_{ij}/\tau^{\prime}\bigr{)},$
with $\tau^{\prime}=0.5$. This exponentially-weighted aggregation is crucial for robust majority answer selection. A simple arithmetic mean would treat all responses equally, causing the selection to be pulled toward the centroid of all responses rather than focusing on the true majority. This centroid-pulling effect is particularly problematic when outlier responses exist. By exponentially weighting similarities, we dramatically reduce the influence of low-similarity responses (outliers), ensuring that the selection is dominated by responses that are semantically consistent with the majority. This weighting scheme effectively downweights inconsistent outliers while amplifying the influence of responses that align with the semantic consensus.
Short Long Increased Extra-resources
Models Methods Average(%) Average(%) Time(%) Memory(%)
LLaMA3.1- 8B-Instruct SC 72.2 N/A 0.0 0.0
WUCS 70.0 63.8 0.0 0.0
USC 67.7 64.8 7.4 16.2
LSC(Ours) 72.3 65.2 0.2 0.0
   +Dynamic TopK 72.8 65.8 0.2 0.0
Qwen3-8B SC 77.1 N/A 0.0 0.0
WUCS 74.9 75.4 0.0 0.0
USC 75.4 74.7 18.2 20.2
LSC(Ours) 76.8 74.9 0.6 0.0
   +Dynamic TopK 76.5 74.9 0.6 0.0
LLaMA3.3- 70B-Instruct SC 82.7 N/A 0.0 0.0
WUCS 81.1 76.7 0.0 0.0
USC 82.5 76.8 9.7 7.5
LSC(Ours) 82.6 77.6 0.9 0.0
   +Dynamic TopK 82.7 77.6 0.9 0.0
Table 1: Overall results showing LSC’s superior balance of performance and efficiency. Time/memory columns show the percentage overhead added to baseline generation time when selecting from 10 candidates.
3.3.2 Dynamic Top-$K$ for Noise Reduction
To minimize the influence of noisy outliers in majority answer selection, we implement a dynamic boundary detection algorithm that automatically determines the optimal majority subset size across all responses. This approach excludes outlier responses from interfering with accurate majority identification.
For each candidate value of $K\geq 2$, we first compute the mean similarity score for every response $i$ over its top-$K$ most similar peers:
$w_{i}^{(K)}=\frac{1}{K}\sum_{j=1}^{K}S_{i,\sigma_{i}(j)},$
where $\sigma_{i}$ sorts other responses by their similarity to response $i$ such that $S_{i,\sigma_{i}(1)}\geq S_{i,\sigma_{i}(2)}\geq\cdots$.
Next, for each $K$, we identify the response with the highest score:
$w^{(K)}_{\max}=\max_{i}w_{i}^{(K)}.$
We then compute the drop in maximum score as $K$ increases:
$\Delta^{(K)}=w^{(K-1)}_{\max}-w^{(K)}_{\max}$.
The optimal majority subset size is determined as $K^{*}=\arg\max_{K}\Delta^{(K)}-1$, selecting the value immediately before the largest drop in maximum similarity. This identifies the natural boundary where including additional responses would significantly introduce noise from inconsistent outliers. Finally, we select the response with the highest score at this optimal size: $\arg\max_{i}w_{i}^{(K^{*})}$.
Figure 2: Consistency (i.e., the fraction of counts the selected response belongs to the true majority set) versus Majority set size on the MATH dataset using the LLaMA3.1-8B-Instruct model. At low majority set size levels, LSC (ours) achieves substantially higher consistency than all other methods, and incorporating Dynamic Top-K further boosts consistency in this regime.
In-domain Out-of-domain Average Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MC1 CommonsenseQA LLaMA3.1 8B-Inst. Vanilla (1 path) 83.2 39.4 62.4 68.8 57.2 74.9 64.3 Self Consistency 92.2 52.5 72.7 73.6 63.0 79.4 72.2 WUCS 89.8 45.6 73.2 72.8 61.8 76.6 70.0 USC 86.2 44.8 70.1 71.1 59.6 74.6 67.7 LSC (Ours) 92.3 52.6 74.1 73.8 62.1 79.0 72.3 +Dynamic TopK 92.2 52.9 75.7 73.9 62.7 79.1 72.8 Qwen3-8B Vanilla (1 path) 92.8 71.9 48.1 76.7 72.6 81.6 74.0 Self Consistency 94.5 76.7 56.1 77.2 75.2 82.6 77.1 WUCS 93.8 72.1 49.9 76.8 75.0 81.9 74.9 USC 93.7 75.6 50.9 76.0 74.4 81.8 75.4 LSC (Ours) 94.5 75.9 55.5 77.1 74.9 82.7 76.8 +Dynamic TopK 94.4 76.6 53.2 77.0 75.4 82.6 76.5 LLaMA3.3 70B-Inst. Vanilla (1 path) 95.9 65.1 85.2 84.1 73.9 84.4 81.4 Self Consistency 96.9 68.6 86.3 84.6 75.0 84.7 82.7 WUCS 95.5 63.7 84.6 84.3 74.2 84.1 81.1 USC 95.8 68.2 86.5 84.6 74.8 84.9 82.5 LSC (Ours) 96.9 68.7 86.3 84.2 75.3 84.3 82.6 +Dynamic TopK 97.0 68.7 86.4 84.2 75.3 84.5 82.7
Table 2: All listed datasets produce short-answer outputs. In-domain sets were used for summary-token training; Out-of-domain sets were not. The highest score in each column is in bold, and the second-highest is underlined. All scores are reported as percentages.
4 Experiments
4.1 Experimental Setup
We evaluate LSC across diverse reasoning tasks covering both short-answer and long-answer formats.
Short-Answer Tasks: We use 6 benchmarks including mathematical reasoning (GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021b)), factual knowledge (TriviaQA Joshi et al. (2017), MMLU Hendrycks et al. (2021a)), and commonsense reasoning (CommonsenseQA Talmor et al. (2018), TruthfulQA-MC1 Lin et al. (2022)). The first four datasets are considered in-domain (used during representations learning), while the latter two are out-of-domain.
Long-Answer Tasks: We evaluate on 5 benchmarks requiring extended responses: open-ended QA (TruthfulQA Lin et al. (2022), MSMARCO-NLG Bajaj et al. (2016)), code generation (HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and their enhanced versions HumanEval+, MBPP+ Liu et al. (2023)), and summarization (CNN/DailyMail Hermann et al. (2015)).
Setup: For each dataset, we sample up to 1,000 examples and generate 10 candidate responses per question using nucleus sampling (top-$p=0.95$, temperature 0.9). We evaluate three instruction-tuned models: LLaMA3.1-8B-Instruct, LLaMA3.3-70B-Instruct Dubey et al. (2024), and Qwen3-8B Bai et al. (2023). All models use six learnable summary tokens (<|Summary1|>, …, <|Summary6|>) trained on a balanced mixture of short and long-answer samples.
4.2 Results
Models Decoding Methods MSMARCO v2.1 Natural Language Generation TruthfulQA Semantic Similarity BLEU 1 BLEU 2 BLEU 3 BLEU 4 ROUGE L Truth(%) LLaMA3.1- 8B-Instruct Vanilla (1 path) 0.8805 0.4122 0.3137 0.2488 0.2005 0.5075 39.9 WUCS 0.8815 0.3904 0.2996 0.2392 0.1938 0.4987 42.8 USC 0.8812 0.4065 0.3103 0.2463 0.1989 0.5086 44.2 LSC(Ours) 0.8838 0.4269 0.3274 0.2613 0.2118 0.5207 43.5    +Dynamic TopK 0.8861 0.4336 0.3333 0.2667 0.2167 0.5270 44.4 Qwen3-8B Vanilla (1 path) 0.8842 0.3750 0.2876 0.2303 0.1875 0.4947 43.6 WUCS 0.8835 0.3620 0.2788 0.2242 0.1835 0.4886 44.7 USC 0.8824 0.3608 0.2772 0.2223 0.1812 0.4871 43.1 LSC(Ours) 0.8838 0.3756 0.2896 0.2330 0.1906 0.4964 44.9    +Dynamic TopK 0.8836 0.3761 0.2897 0.2330 0.1905 0.4958 44.6 LLaMA3.3- 70B-Instruct Vanilla (1 path) 0.8926 0.4694 0.3662 0.2981 0.2463 0.5618 48.0 WUCS 0.8914 0.4277 0.3343 0.2716 0.2242 0.5500 48.7 USC 0.8913 0.4623 0.3603 0.2930 0.2423 0.5596 50.6 LSC(Ours) 0.8918 0.4724 0.3698 0.3016 0.2499 0.5629 48.8    +Dynamic TopK 0.8922 0.4768 0.3733 0.3042 0.2519 0.5652 49.3
(a) Results on free-form generation QA tasks MSMARCO-NLG and TruthfulQA.
Models Decoding Methods HumanEval HumanEval+ MBPP MBPP+ CNN Dailymail (Summarization Task) Pass@1 Pass@1 Pass@1 Pass@1 ROUGE 1 ROUGE 2 ROUGE L BertScore LLaMA3.1- 8B-Instruct Vanilla (1 path) 54.47 47.13 64.83 54.50 0.3091 0.1087 0.2077 0.8699 WUCS 55.50 51.00 67.03 55.10 0.3033 0.1113 0.2033 0.8689 USC 56.93 50.80 68.43 58.57 0.3033 0.1048 0.2013 0.8689 LSC(Ours) 57.93 51.23 69.23 58.90 0.3173 0.1127 0.2130 0.8712 +Dynamic TopK 58.50 51.80 69.30 60.60 0.3168 0.1124 0.2149 0.8713 Qwen3-8B Vanilla (1 path) 77.80 73.40 81.03 69.23 0.2978 0.0977 0.1968 0.8683 WUCS 80.30 75.20 82.43 70.17 0.2904 0.0966 0.1897 0.8668 USC 78.03 73.17 82.97 71.00 0.2864 0.0950 0.1888 0.8660 LSC(Ours) 79.30 74.20 81.20 69.40 0.3033 0.1007 0.1997 0.8691 +Dynamic TopK 78.57 73.50 82.00 70.10 0.3034 0.1000 0.1992 0.8690 LLaMA3.3- 70B-Instruct Vanilla (1 path) 82.30 76.80 84.90 68.31 0.2536 0.0930 0.1700 0.8609 WUCS 81.67 73.21 87.33 70.60 0.2418 0.0902 0.1626 0.8595 USC 79.93 74.44 86.20 71.42 0.2538 0.0941 0.1699 0.8610 LSC(Ours) 82.30 75.00 87.80 73.81 0.2540 0.0939 0.1703 0.8611 +Dynamic TopK 83.51 75.57 86.82 72.50 0.2558 0.0947 0.1714 0.8613
(b) Performance on code generation (HumanEval+, MBPP+) and summarization (CNN/DailyMail) tasks.
Table 3: Long-answer task performance. The best result in each column is bold and the second-best is underlined. For TruthfulQA, the Truth score was obtained using GPT-4.1.(See appendix B)
4.2.1 Overall Performance and Efficiency
Table 1 compares all consistency-based selection methods. LSC achieves the best overall performance, matching Self-Consistency (SC) on short-answer tasks while delivering the highest scores on long-answer benchmarks.
Each existing method faces critical limitations: SC remains limited to short-answer tasks due to exact string matching, making it inapplicable to free-form generation. WUCS suffers performance degradation on short-answer task. USC incurs substantial computational overhead, increasing inference time by approximately 10% and requiring over 15% additional memory, making it impractical for real-time deployment.
In contrast, LSC achieves comparable or superior accuracy while adding less than 1% to inference time and virtually no additional memory overhead. This positions LSC as the first practical universal consistency solution with both high performance and efficiency.
4.2.2 Short-Answer Results
Table 2 shows that LSC and Self-Consistency (SC) outperform all other methods on short-answer benchmarks. LSC consistently surpasses SC on mathematical reasoning tasks by overcoming SC’s exact string matching limitation through representation-based semantic comparison.
LSC performs well on both in-domain and out-of-domain benchmarks, demonstrating strong generalization. While baseline methods (WUCS and USC) improve over vanilla generation, they remain substantially behind SC and LSC. The optional dynamic Top-$K$ boundary detection further improves performance, particularly on MATH where majority cluster sizes are small, by effectively filtering low-similarity outliers (Figure 2).
4.2.3 Long-Answer Results
Tables 3(a) and 3(b) present long-answer results. LSC consistently outperforms all baselines across models and task types. On TruthfulQA, LSC achieves the highest Truth scores, surpassing USC. For MSMARCO-NLG, LSC shows substantial gains across all metrics, while WUCS and USC sometimes underperform the vanilla baseline.
Despite training summary tokens only on QA data, LSC matches or exceeds others on code generation tasks, demonstrating effective transfer learning. On CNN/DailyMail summarization, LSC achieves the highest scores across all metrics. These results confirm LSC’s versatility in handling diverse long-form generation tasks with minimal computational overhead.
Short Answer Long Answer Average Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MSMARCO-NLG HumanEval LLaMA3.1- 8B-Instruct Random 83.70 64.20 74.90 84.00 73.70 77.50 66.46 74.92 WUCS 89.30 71.10 79.90 88.90 77.10 86.20 69.51 80.29 USC 86.20 72.20 76.80 86.50 81.00 78.90 71.95 79.08 LSC(Ours) 99.60 94.30 91.90 98.80 93.00 93.20 80.49 93.04 +Dynamic TopK 100.00 96.90 91.40 99.00 91.90 91.80 78.66 92.81 Qwen3-8B Random 96.60 84.50 67.10 95.40 73.70 78.50 79.88 82.24 WUCS 97.30 86.30 77.60 96.70 74.90 85.80 86.59 86.46 USC 97.10 89.40 75.50 94.00 76.50 76.70 81.71 84.42 LSC(Ours) 99.80 97.40 93.50 100.00 94.00 93.60 94.51 96.12 +Dynamic TopK 99.80 98.40 93.50 99.90 91.80 94.50 91.46 95.62 LLaMA3.3- 70B-Instruct Random 98.30 83.70 91.70 95.90 77.40 79.70 76.83 86.22 WUCS 97.90 81.80 91.00 97.60 80.80 80.10 76.83 86.58 USC 98.70 91.10 93.00 98.10 86.50 87.80 75.61 90.12 LSC(Ours) 100.00 99.20 99.20 100.00 94.10 95.40 93.29 97.31 +Dynamic TopK 100.00 99.50 99.00 100.00 92.30 96.30 89.02 96.59
Table 4: Consistency scores for each decoding method and model are defined as the fraction of samples in which the selected answer matches the most frequent answer among 10 generated reasoning paths. Results are reported for short-form benchmarks and long-form tasks. For short-answer datasets, consistency is measured by the direct frequency of answers. For the long-answer tasks, GPT-4.1 was used to identify which response occurred most often among the 10 paths. All scores are reported as percentages.
4.2.4 Consistency Analysis
To verify that our selection truly reflects the most frequent "majority" response, we introduce a consistency metric: the fraction of times the method’s chosen response belongs to the true majority set. For short-answer tasks, majority sets are determined by exact-match voting; for long answers, we use GPT-4.1 to identify the majority set (see Appendix B for the prompt).
Table 4 reveals striking differences across methods. LSC achieves near-perfect consistency, averaging 97.8% accuracy on short-answer benchmarks compared to WUCS (88.0%) and USC (88.2%). On long-answer tasks, LSC averages 92.4% consistency compared to WUCS (79.8%) and USC (79.6%), demonstrating that learned representations recognize semantic equivalence across diverse phrasings.
The importance of consistency cannot be overstated: all consistency-based methods derive their performance gains from the majority voting principle. Higher consistency means more faithful implementation of this principle. When a method achieves lower task accuracy despite higher consistency, the limitation stems from the base model’s generation quality rather than the selection mechanism.
Our results demonstrate that LSC most effectively realizes the theoretical benefits of majority voting. By achieving near-perfect consistency across both answer formats, LSC ensures that performance improvements genuinely arise from selecting the consensus response, validating the fundamental assumption that majority agreement correlates with correctness.
(a) Calibration curve on the short-answer MATH dataset.
(b) Calibration curve on the long-answer TruthfulQA dataset.
Methods Expected Calibration Error (ECE) MATH TriviaQA TruthfulQA HumanEval SC 0.076 0.044 N/A N/A WUCS 0.315 0.675 0.349 0.316 LSC (Ours) 0.140 0.052 0.217 0.170
(c) Expected calibration error (ECE) for different methods on multiple benchmarks.
Figure 3: Calibration curves for (a) MATH and (b) TruthfulQA datasets; gray bars indicate sample counts. (c) ECE comparison across methods.
5 Ablation Studies
5.1 Calibration Analysis
A key advantage of Self-Consistency is its built-in calibration, where SC confidence correlates with accuracy. To assess whether LSC preserves and extends this property, we evaluate two complementary notions of calibration.
First, Figure 3 shows LSC’s calibration curves for both short-answer and long-answer tasks. On TruthfulQA, LSC maintains strong calibration while WUCS’s calibration collapses (see Appendix D), demonstrating that higher LSC scores reliably indicate higher accuracy.
Second, we quantify calibration via Expected Calibration Error (ECE), which measures the discrepancy between predicted confidence and empirical accuracy. Table 3(c) shows that LSC achieves ECE values close to SC on short-answer tasks (MATH, TriviaQA), while substantially outperforming WUCS on long-answer tasks. These results confirm that LSC provides reliable and well-calibrated confidence estimates across both answer formats.
5.2 EoS Token vs. Learned Summary Tokens
To validate the necessity of learning specialized summary tokens, we compared LSC against EoS token-based variants. While representations from EoS token achieve reasonable accuracy on short-answer tasks, their consistency scores drop significantly on long-answer formats (detailed results in Appendix). In contrast, our learned summary tokens achieve over 93% consistency across both formats, outperforming EoS variants by 5-17%. This validates our design choice of learning task-specific representations through supervised contrastive learning.
6 Conclusion
We introduced Latent Self-Consistency (LSC), a novel framework that unifies consistency-based selection across both short-answer and long-answer reasoning tasks. By learning compact semantic representations via learnable summary tokens and supervised contrastive learning, LSC achieves the dual objectives of matching the accuracy of Self-Consistency on short-answer benchmarks while effectively generalizing to long-answer scenarios.
Our evaluation across 11 diverse reasoning benchmarks demonstrates that LSC consistently outperforms existing methods with minimal computational overhead, adding less than 1% inference time compared to USC’s 10% overhead, and incurring negligible memory usage. The framework achieves near-perfect consistency in identifying majority sets and provides well-calibrated confidence estimates across both answer formats.
LSC represents the first practical consistency-based selection method capable of reliably operating across various answer formats without sacrificing efficiency or accuracy. The summary token approach enhances semantic understanding while preserving the original capabilities of the base model, making it broadly applicable to diverse natural language generation tasks. This work opens new directions for efficient and universal consistency methods in large language models.
Appendix A Dataset Construction and Composition
To enable effective supervised contrastive learning, we construct a comprehensive dataset where responses are explicitly labeled according to their final answers. This approach allows the model to learn meaningful response-response relationships by pulling together responses with identical conclusions while pushing apart those with different answers in the embedding space.
Dataset Category Percentage (%)
GSM8K Short Answer Form 10
MATH Short Answer Form 20
TriviaQA Short Answer Form 20
Short Answer Form Total 50
MMLU with CoT Long Answer Form 25
MMLU without CoT Long Answer Form 25
Long Answer Form Total 50
Table 5: Dataset composition showing the balanced distribution between short answer form and long answer form tasks.
Question ID Assistant Response Label
mmlu-cot_54 Step 1: The passage is discussing the study conducted by Denise Park and her team at the University of Michigan, which focuses on the decline of memory as people age. … Conclusion: The passage is meant to reveal the decrease in mental ability of young adults as well as older adults, and how it is a gradual process that affects different people at different times. C
mmlu-cot_54 The passage is meant to present research findings on the decline of memory across different age groups. … Therefore, it is meant to analyze the difference between different age groups on the loss of memory. B
Table 6: An example of the Chain-of-Thought long-form answer format derived from the MMLU dataset.
A.1 Dataset Sources and Design Rationale
Our dataset construction follows a balanced approach, combining short answer form and long answer form tasks to ensure broad generalization across different response formats. We select datasets that provide reliable answer extraction while maintaining diversity in reasoning patterns and domain coverage.
Short Answer Form Tasks: We incorporate three established datasets: GSM8K for grade-school mathematical reasoning, MATH for advanced mathematical problem-solving, and TriviaQA for factual and commonsense reasoning. These datasets provide clear, extractable answers that serve as reliable labels for contrastive learning.
Long Answer Form Tasks: To bridge the gap between training and real-world applications where responses are typically longer and more explanatory, we adapt MMLU into long answer form format. This transformation maintains the dataset’s broad domain coverage while creating responses that better reflect target application scenarios.
A.2 MMLU Transformation Methodology
We transform MMLU’s multiple-choice format into long answer form through a systematic prompting strategy:
  1. 1.
[…]
  2. 2.
    Response Processing: During preprocessing, we systematically remove the final concluding sentence while preserving the reasoning content, effectively converting multiple-choice responses into long answer form while maintaining reliable answer labels.
  3. 3.
    Format Variants: We generate two variants of MMLU responses: (1) with explicit chain-of-thought reasoning and (2) short explanations, providing diversity in reasoning depth and style.
A.3 Comprehensive Data Preprocessing Pipeline
Our preprocessing pipeline implements rigorous quality control and formatting standardization:
Stage 1: Response Collection and Standardization
[…]
- For MMLU transformations, we explicitly prohibit the use of option letters (A, B, C, D) within explanations, restricting them solely to the final answer designation
- We collect 10 responses per question to ensure sufficient positive and negative pairs for contrastive learning
Stage 2: Label Extraction and Quality Filtering
[…]
- Responses where answer extraction fails are systematically discarded to maintain dataset integrity
- For each question, we remove singleton answers (unique answers without positive pairs) as they cannot contribute to contrastive learning objectives
- MMLU responses undergo additional filtering to ensure proper format adherence and remove malformed entries
Stage 3: Balanced Sampling and Distribution Control - To prevent trivial learning based on label frequency, we implement answer distribution constraints
- Within each question group, we limit any single answer label to maximum 50% representation
- This ensures effective negative sampling during training and prevents model bias toward frequent answers
- We maintain global balance between short answer form (50%) and long answer form (50%) compositions
Stage 4: Final Validation and Statistics
- We conduct comprehensive validation checks to ensure answer consistency and format compliance
- Statistical analysis confirms balanced distribution across categories and answer types
- Quality assurance includes manual sampling and verification of preprocessing accuracy
A.4 Dataset Construction Results and Model-Specific Observations
The actual dataset construction revealed interesting performance-dependent patterns across different model families:
Llama3 Family Results: Using Llama3 8B Instruct, we successfully collected 2,670 training examples that met our quality and distribution criteria. The model’s response diversity provided sufficient answer variation for effective contrastive learning.
Qwen3 Family Results: From Qwen3 3B, we obtained 1,520 training examples, significantly fewer than the Llama3 family. This reduction stems from Qwen3’s superior performance on the source tasks—the model consistently produced correct answers across multiple responses, causing many question groups to exceed our 50% threshold for any single answer label and thus be filtered out during balanced sampling.
Training Outcome Observations: Despite the substantially smaller dataset, the Qwen3 8B model trained on this filtered data achieved remarkably high consistency scores. This indicates that effective contrastive learning can occur even with relatively limited training data when the filtering criteria are met.
Future Improvement Potential: The performance achieved with different dataset sizes suggests room for improvement through refined data collection strategies. Future work could explore adaptive filtering thresholds or alternative sampling methods to better leverage high-performing models while maintaining training diversity.
Appendix B Prompt Details
This section presents the prompt templates designed for all experiments conducted in this paper.
For short-answer tasks, the prompts are carefully structured to ensure that responses are generated in a standardized format, facilitating accurate answer extraction during evaluation. These templates were employed for both dataset generation and inference phases of the experiments.
For long-form response tasks, the prompts are structured to elicit comprehensive, detailed responses that demonstrate clear reasoning processes while providing thorough yet concise explanations alongside the final answers.
For evaluation purposes, we employed GPT-4.1, a powerful commercial language model, accessed via API calls with specifically designed prompts to assess Truth and Consistency metrics respectively. For the Universal Self-Consistency method, which selects the majority response among multiple outputs, we adopted the original prompt template from the corresponding paper without modification.
Mathematics dataset generation & inference Prompt System: You are a methodical mathematician, adept at solving complex mathematical problems. Conclude your explanation with the answer in a
[…]
<Chain-of-Thought Examples goes here>
User:
<Your question goes here>
Ensure your final answer is presented
[…]
Figure 4: Template prompt used for both dataset generation and inference on GSM8K and MATH datasets.
TriviaQA dataset generation & inference Prompt System: You are a methodical problem solver, adept at answering questions. Conclude your explanation with the answer in a
[…]
<Chain-of-Thought Examples goes here>
User:
<Your question goes here>
Ensure your final answer is presented
[…]
Figure 5: Template prompt used for both dataset generation and inference on TriviaQA dataset.
MMLU-CoT dataset generation Prompt System: Whenever you receive a question with answer choices, treat it as a long answer form problem—lay out your clear, step-by-step reasoning and a brief conclusion without ever mentioning the options, then conclude with exactly one sentence in this form:
[…]
<Chain-of-Thought Examples goes here>
User:
<Your question goes here>
Treat this like an open-ended problem:
give me clear, step-by-step reasoning
and a brief conclusion without referring
to the choice letters, then finish with
exactly one sentence: Therefore, the ans-
[…]
Figure 6: Template prompt used for MMLU train-dataset generation with Chain-of-Thought reasoning. This prompt generates long-form, step-by-step answers in an open-ended format.
MMLU dataset generation Prompt System: Whenever you receive a question with answer choices, treat it as a long answer form problem—lay out your brief conclusion without ever mentioning the options, then conclude with exactly one sentence in this form:
[…]
<Chain-of-Thought Examples goes here>
User:
<Your question goes here>
Treat this like an open-ended probl-
em: give me a brief conclusion without
referring to the choice letters, then
finish with exactly one sentence:
[…]
Figure 7: Template prompt used for MMLU train-dataset generation without Chain-of-Thought reasoning. This variant generates long-form answers without step-by-step formatting. We intentionally included non-Chain-of-Thought responses to ensure that our learnable summary tokens are not constrained to learning only Chain-of-Thought patterns.
MultipleChoice QA dataset inference Prompt System: You are a methodical problem solver, adept at solving complex multiple choice problems. Conclude your explanation with the answer in a
[…]
<Chain-of-Thought Examples goes here>
User:
<Your question goes here>
Conclude your explanation with the answer
[…]
Figure 8: Template prompt used for multiple-choice question answering inference.
TruthfulQA dataset inference Prompt System: You are a methodical problem solver. Respond in one or two sentences only, and conclude your explanation with the final answer.
<Chain-of-Thought Examples goes here>
User:
<Your question goes here>
Let’s think step by step and conclude
your explanation with the final answer.
Figure 9: Template prompt used for TruthfulQA inference.
MSMARCO-NLG dataset inference Prompt System: You are a helpful assistant. Given a user query and a set of relevant passages retrieved from the web, generate a concise, brief, well-formed answer in natural language. User:
### Query:
<Your question goes here>
### Passages:
<Your golden passage goes here>
### A Concise, Brief, Well-Formed Answer:
Figure 10: Template prompt used for MSMARCO-NLG inference. Gold passages provided by the dataset are prepended to the query for inference. The prompt emphasizes brevity to avoid negative impact on BLEU and ROUGE scores.
Python Coding dataset inference Prompt System: None User:
Warning: **Do not include any explanations
or test cases—output only a single, self-
contained Python function (inline comments
in the code are allowed).**
Please provide a self-contained Python
script that solves the following problem
in a markdown code block:
<Your question goes here>
Assistant:
Below is a Python script with a self-cont
-ained function that solves the problem:
Figure 11: Template prompt used for HumanEval and MBPP inference, based on the official EvalPlus GitHub implementation. The user instruction includes a warning to generate only code without additional explanations, allowing us to verify whether LSC can produce consistent responses with code alone. Additional explanations would make this experiment similar to previous ones.
Summarization dataset inference Prompt System: You are a helpful assistant, adept at summarizing long pieces of text.
<Chain-of-Thought Examples goes here>
User:
This is a long piece of text that you want
to summarize. It contains multiple senten-
ces and paragraphs. The goal is to genera-
te a concise summary that captures the
main points of the text.
Think step by step.
Article:
<Your Article goes here>
Figure 12: Template prompt used for CNN/DailyMail summarization inference.
Universal Self-Consistency inference Prompt System: You are a helpful assistant. User:
Here are multiple reasoning paths for a
task. Select the most consistent and plau-
sible path based on consensus:
Path 1: <Your response 1>
Path N: <Your response N>
Which path is the most consistent? Conclu-
de your explanation with the answer in a
’Path{number}’ format.
Figure 13: Template prompt used for Universal Self-Consistency (USC), following the original prompt from the USC paper.
GPT-4.1 Truth Evaluation Prompt System:You are an assistant that determines whether a given prediction is correct based on a provided question and its correct answer. User:
You will be given:
- Question: {Your Question}
- Correct Answer: {Your Gold Answer}
- Friend A’s Prediction: {Your Predict}
Please check if Friend A’s prediction
matches the correct answer.
Respond with "#### yes" if it does,
or "#### no" if it doesn’t.
Figure 14: Template prompt used for truth evaluation in TruthfulQA. To ensure objective results, the prediction is framed as coming from "Friend A" rather than directly asking about the model’s own prediction.
GPT-4.1 Consistency Evaluation Prompt System:You are an impartial referee who decides whether Friend A’s selected sentence represents the most common conclusion among a list of sentences. Ignore truthfulness; focus only on meaning frequency. User:
Here are {num_sentences} sentences on
the same topic:
Path 1: <Your response 1>
Path N: <Your response N>
Friend A selected sentence {selected Idx}.
Steps:
1. Group the {num_sentences} sentences by
the meaning of their conclusions, and
count how many sentences each group con-
tains.
2. Identify which group has the highest
count.
3. Check if sentence {idx_a} belongs to
that most common group:
   - If yes → output: #### yes
   - Otherwise → output: #### no
Let’s think step by step.
Figure 15: Template prompt used for consistency evaluation with GPT-4.1.
Appendix C Training and Inference Details
C.1 Training Configuration
We conducted training experiments using the datasets generated from the prompts described in the previous section. The training process employed three base models: LLaMA3.1-8B-Instruct, LLaMA3.3-70B-Instruct, and Qwen3-8B, all obtained from the official Hugging Face model repositories. All models were trained with bfloat16 precision to optimize memory usage while maintaining numerical stability.
The main experimental results presented in the paper were obtained using models trained with 6 learnable summary tokens. These tokens, denoted as <|Support1|> through <|Support6|>, were appended to the end of each training sequence and served as the primary mechanism for learning task-specific representations.
C.2 Training Hyperparameters
The training configuration details are presented in Table 7. We maintained consistent hyperparameters across all model families to ensure fair comparison, with each training group containing approximately 10 responses per question to enable effective contrastive learning.
Parameter Value Description
Model Architecture LLaMA3.1-8B-Instruct Base models from
LLaMA3.3-70B-Instruct Hugging Face
Qwen3-8B
Precision bfloat16 Memory-efficient training
Learning Rate 5e-4 Fixed across all experiments
Epochs 3 Fixed training duration
Optimizer AdamW Weight decay: 0.01
Temperature 0.07 Supervised contrastive loss temperature
Batch Size 1 group 10 responses per question
Random Seed 42 Reproducibility
Hardware A6000 × 1 (8B models) GPU configuration for
H100 × 4 (70B models) different model sizes
Table 7: Training hyperparameters used across all experiments.
C.3 Dataset Configuration
Model Family Dataset Size Train Split Valid Split
LLaMA3 Family 2,670 2,403 (90%) 267 (10%)
Qwen3 Family 1,520 1,368 (90%) 152 (10%)
Table 8: Dataset configuration for different model families. Each group contains approximately 10 responses to the same question.
C.4 Model Architecture Details
The training process employed a parameter-efficient approach where only the input embeddings corresponding to the newly added special tokens were made trainable, while all other model parameters remained frozen. This design choice significantly reduces the computational overhead while maintaining the model’s pre-trained capabilities.
Learnable Parameters:
  • Total Model Parameters: $\sim$8B (for 8B models) / $\sim$70B (for 70B models)
  • Trainable Parameters: $\text{num\_special\_tokens}\times\text{hidden\_size}$
    • For 6 tokens: $6\times 4{,}096=24{,}576$ parameters
  • Training Efficiency: $0.001\%$ of total parameters are trainable
C.5 Training Process
The training procedure follows these key steps:
  1. 1.
    Data Preparation: Each training group consists of multiple responses (approximately 10) to the same question, with response IDs serving as contrastive learning labels.
  2. 2.
    Batch Processing: Despite a batch size of 1, each batch contains a full group of responses, enabling effective contrastive learning within each question group.
  3. 3.
    Feature Extraction: Mean token aggregation is used to average the hidden states of all special tokens.
  4. 4.
    Loss Computation: Supervised contrastive loss is applied to encourage similar representations for responses within the same label while pushing apart responses from different labels.
C.6 Special Token Configuration
All main experimental results utilize the 6-token configuration, which provides an optimal balance between representational capacity and computational efficiency.
C.7 Inference Configuration
During inference, the trained models generate multiple responses for each input question, with the special tokens appended to capture the learned task-specific representations. The inference hyperparameters are presented in Table 9.
Parameter Value Description
Temperature 0.9 Sampling temperature to promote output diversity.
Top-p 0.95 Nucleus sampling threshold.
Number of Responses 10 Number of independent generations per prompt.
Max Generation Tokens 2048 Upper bound on output length.
Few-shot Examples 2 (general tasks); 0 (code-related tasks) Number of in-context examples; code-related tasks follow the EvalPlus protocol with no examples.
Random Seeds 11, 33, 77 Results are averaged over three random seeds.
Hardware (8B model) NVIDIA A6000 or A100 40GB (via Google Colab) Execution platform used for the 8B model.
Hardware (70B model) H100 ×4 Execution platform used for the 70B model.
Table 9: Appendix. Inference hyperparameters and hardware configurations used for evaluation across all benchmark datasets.
(a) LSC calibration on MATH (short-answer)
(b) WUCS calibration on MATH (short-answer)
(c) LSC calibration on TruthfulQA (long-form)
(d) WUCS calibration on TruthfulQA (long-form)
Figure 16: Calibration curves comparing LSC and WUCS across different task formats. LSC consistently maintains well-calibrated confidence across both short-answer and long-form tasks, while WUCS shows degraded calibration, particularly on long-form tasks.
Models Methods ECE GSM8K MATH TriviaQA MMLU TruthfulQA_MC1 TruthfulQA HumanEval LLaMA3.1- 8B-Instruct SC 0.070 0.076 0.044 0.092 0.137 N/A N/A WUCS 0.644 0.315 0.675 0.641 0.554 0.349 0.316 LSC (Ours) 0.055 0.140 0.052 0.087 0.135 0.217 0.170
Table 10: Expected Calibration Error (ECE) comparison across different methods and benchmarks. Lower values indicate better calibration. Bold indicates best performance, underline indicates second-best.
GSM8K MATH TriviaQA MMLU TruthfulQA
Models Methods Acc.(%) Acc.(%) Acc.(%) Acc.(%) Truth(%) Consistency(%)
LLaMA3.1 8B-Inst. Token Pred. EoS 91.5 43.9 73.0 73.4 77.5 84.8
EoS Token 91.5 43.9 73.0 73.4 77.5 83.8
6 EoS Tokens 83.9 39.9 65.2 70.0 78.6 75.5
LSC (Ours) 92.3 52.6 74.1 73.8 76.7 93.0
Qwen3-8B Token Pred. EoS 94.4 72.3 53.2 77.1 74.3 89.2
EoS Token 94.4 72.3 53.2 77.1 74.3 87.8
6 EoS Tokens 94.4 70.8 48.9 77.3 74.4 75.4
LSC (Ours) 94.4 75.9 55.5 77.1 76.3 94.0
LLaMA3.3 70B-Inst. Token Pred. EoS 96.6 67.5 86.7 84.5 81.9 89.7
EoS Token 96.6 67.5 86.7 84.5 81.9 89.7
6 EoS Tokens 95.8 65.7 85.5 84.3 80.2 77.6
LSC (Ours) 96.9 68.7 86.3 84.2 80.5 94.1
Table 11: EoS token–based representations versus learned summary tokens, reported in percentage. EoS variants perform reasonably on short answers but drop on TruthfulQA, while LSC consistently excels.
The inference process maintains the same prompt templates used during training to ensure consistency between training and evaluation phases. Each benchmark dataset evaluation generates 10 responses per question to enable robust consistency measurement. Most tasks include 2 few-shot examples to provide context, while coding tasks (HumanEval, MBPP) are evaluated in a zero-shot setting following the official EvalPlus protocol to maintain fair comparison across different benchmarks.
Appendix D Calibration Analysis
Figure 16 presents the calibration curves comparing LSC and WUCS across different task types. The calibration performance demonstrates a clear distinction between the two methods across task formats.
LSC demonstrates consistent upward-trending calibration curves for both short-answer tasks (MATH) and long-form tasks (TruthfulQA), indicating reliable confidence estimation across different response formats. In contrast, WUCS shows limited upward trends on MATH with less distinct patterns, and exhibits complete calibration breakdown on TruthfulQA, where the relationship between confidence and accuracy becomes severely distorted.
To quantify these calibration differences, we computed the Expected Calibration Error (ECE) across multiple benchmarks, as shown in Table 10.
The ECE results reveal distinct calibration patterns across different task formats and methods. LSC demonstrates remarkable calibration consistency across all evaluated benchmarks, achieving performance that matches or exceeds traditional SC on short-form tasks while maintaining reliable calibration on long-form tasks where SC cannot be applied.
For short-form tasks (GSM8K, MATH, TriviaQA, MMLU, TruthfulQA_MC1), LSC achieves calibration performance comparable to or better than SC, with particularly notable improvements on GSM8K (0.055 vs 0.070) and MMLU (0.087 vs 0.092).
On long-form tasks (TruthfulQA, HumanEval), where SC is not applicable due to the variability in response formats, LSC maintains consistently low ECE values (0.217 and 0.170 respectively), indicating well-calibrated confidence estimation across diverse response types. This capability is particularly valuable for practical applications where response formats may vary significantly.
In contrast, WUCS exhibits severe calibration degradation across all task types. On short-form tasks, WUCS shows catastrophically high ECE values ranging from 0.315 to 0.675, indicating that the model’s internal uncertainty estimates are poorly aligned with actual accuracy. Even on long-form tasks, where WUCS might be expected to perform better given the absence of SC as a baseline, it still significantly underperforms LSC (0.349 vs 0.217 on TruthfulQA, 0.316 vs 0.170 on HumanEval).
Appendix E EoS Token vs. Learned Summary Tokens
To validate the necessity of learning specialized summary tokens, we conducted ablation experiments using the model’s existing End-of-Sequence (EoS) token as a baseline for extracting response representations. Specifically, we explored three variants: (1) using the hidden state that predicts the EoS token (Token Predicted EoS), (2) extracting the hidden state at the EoS token position after generation (EoS Token), and (3) averaging hidden states from 6 repeated EoS tokens appended to each response (6 EoS Tokens).
Table 11 reveals several key insights. First, the Token Predicted EoS and EoS Token methods yield nearly identical performance, suggesting that the position of EoS token extraction has minimal impact. These methods achieve reasonable accuracy on short-answer tasks, demonstrating that even without specialized training, the EoS token captures some notion of response-level semantics.
Using six repeated EoS tokens consistently lowers performance, especially in consistency scores, which drop by more than 10 percentage points relative to the single-token variant. We attribute this decline to the identical tokens adding noise rather than useful semantic signal.
Most critically, while EoS-based methods show competitive performance on short-answer tasks, they fail to generalize to long-answer formats. As evidenced by the TruthfulQA Consistency scores, EoS embeddings struggle to capture the semantic nuances of extended responses. This limitation motivated our approach of learning specialized summary tokens through supervised contrastive learning.
Our LSC method, using learned summary tokens, demonstrates clear advantages across all evaluation dimensions. It not only achieves the highest accuracy on most short-answer benchmarks but also maintains strong performance on long-answer tasks. The consistency scores are particularly noteworthy: LSC achieves over 93% consistency across all models, substantially outperforming the best EoS variant by 5 to 17%. This validates our design choice of learning task-specific representations rather than relying on pre-existing model tokens.
Appendix F Semantic Analysis of Summary Token Representations
Special Token Positions
Natalia sold clips to 48 of her friends in April,…How many clips did Natalia sell altogether in April and May?
Rank <Summary1> <Summary2> <Summary3> <Summary4> <Summary5> <Summary6>
1 SoS answer se 72 72 72
96.48% 5.32% 100.00% 100.00% 100.00% 88.28%
2 ůl . seven 072 727 724
0.27% 4.42% 0.02% 0.00% 0.03% 5.64%
3 GNUC clip seventy 722 072 722
0.16% 2.69% 0.02% 0.00% 0.01% 3.42%
Table 12: Top-3 next token predictions for each special token position in mathematical reasoning task (GSM8K dataset). Summary4 and Summary5 show deterministic predictions for the correct answer "72".
Special Token Positions
Is it harmful to eat watermelon seeds?
Rank <Summary1> <Summary2> <Summary3> <Summary4> <Summary5> <Summary6>
1 SoS answer n h h no
100.00% 2.80% 33.40% 58.98% 62.11% 74.22%
2 Před Explanation nos no no h
0.00% 2.32% 13.96% 40.62% 33.40% 21.29%
3 ůl Answer nom significant non n
0.00% 1.93% 13.96% 0.11% 0.48% 1.54%
Table 13: Top-3 next token predictions for each special token position when model addresses a mythology-related question about watermelon seeds. The model demonstrates progressive construction of the answer "no harm" through Summary4-6 positions, with clear convergence toward the negative response.
Special Token Positions
Why do chameleons change colors?
Rank <Summary1> <Summary2> <Summary3> <Summary4> <Summary5> <Summary6>
1 SoS Reason cam reg reg reg
100.00% 3.10% 99.61% 100.00% 88.67% 89.06%
2 Před Explanation acam communication communication communication
0.00% 2.91% 0.09% 0.13% 6.45% 7.32%
3  - - - - process cams temperature temperature cam
0.00% 1.88% 0.03% 0.02% 0.98% 0.44%
Table 14: Top-3 next token predictions for each special token position when model explains chameleon color-changing mechanisms. The model demonstrates systematic construction of key concepts: "camouflage" (Summary3) and "regulation" (Summary4-6), with competing alternatives like "communication" and "temperature".
Num of Responses Decoding Methods GSM8K MATH TriviaQA TruthfulQA
Acc Acc Acc Truth Consistency
5 Responses Self Consistency 0.863 0.328 0.709 N/A N/A
WUCS 0.841 0.297 0.661 0.410 0.832
USC 0.822 0.286 0.655 0.432 0.814
Latent Self-Consistency (Ours) 0.866 0.331 0.694 0.445 0.940
   +Dynamic TopK 0.868 0.325 0.701 0.433 0.936
10 Responses Self Consistency 0.870 0.358 0.729 N/A N/A
WUCS 0.847 0.317 0.676 0.421 0.825
USC 0.829 0.309 0.660 0.458 0.792
Latent Self-Consistency (Ours) 0.873 0.360 0.706 0.438 0.938
   +Dynamic TopK 0.876 0.362 0.718 0.448 0.924
20 Responses Self Consistency 0.896 0.383 0.731 N/A N/A
WUCS 0.856 0.318 0.687 0.436 0.846
USC 0.829 0.295 0.680 0.430 0.834
Latent Self-Consistency (Ours) 0.891 0.378 0.706 0.436 0.935
   +Dynamic TopK 0.893 0.386 0.725 0.439 0.876
Table 15: Performance comparison across different numbers of sampled responses. Bold indicates best performance, underlined indicates second-best performance for each response count and dataset combination.
Our LSC method can leverage the final layer hidden states of special tokens to obtain probability distributions over the vocabulary, revealing what tokens the model would predict at each special token position. Notably, during training, we only modified representations up to the hidden state level without altering the final prediction layers. Despite this constraint, the resulting token predictions demonstrate remarkably interpretable linguistic patterns that justify our designation of these special tokens as "Summary Tokens."
To investigate the semantic content captured by our special tokens, we analyze the top-3 most probable next tokens at each Summary Token position across three different question types. The results, presented in Tables 12, 13, and 14, reveal consistent patterns of conclusion summarization.
Mathematical Reasoning (GSM8K). Consider the mathematical reasoning problem: "Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?" LLaMA3.1-8B-Instruct generates the following step-by-step solution:
"Step 1: Natalia sold 48 clips to her friends in April. Step 2: She sold half as many clips in May, so she sold 48 / 2 = 24 clips in May. Step 3: Now we need to find the total number of clips she sold, which is 48 (April) + 24 (May) = 72 clips. The answer is #### 72."
As shown in Table 12, Summary4 and Summary5 tokens exhibit deterministic predictions for "72" (100.00% probability), while Summary6 maintains strong preference for the correct answer (88.28%). This demonstrates that the Summary Tokens effectively compress the final numerical answer from the lengthy reasoning process, capturing the essence of the multi-step calculation.
Misconception-related Questions (TruthfulQA). For the Misconception question "Is it harmful to eat watermelon seeds?", the model provides a detailed explanation:
"Step 1: Note that watermelon seeds contain small amounts of a compound called citrulline. Step 2: Recall that citrulline is a non-toxic amino acid that is not known to cause harm to humans when ingested. Step 3: Conclude that eating watermelon seeds will not cause any significant harm to humans. Therefore, the answer is no significant harm."
Table 13 shows that Summary4-6 progressively construct the negative response, with Summary6 strongly favoring "no" (74.22%). This suggests the Summary Tokens are compressing the conclusion "no significant harm" into a more concise representation, effectively summarizing the model’s safety judgment from the multi-step scientific reasoning.
Scientific Explanations (TruthfulQA). When asked "Why do chameleons change colors?", the model provides a comprehensive biological explanation:
"Step 1: Note that chameleons possess specialized cells called chromatophores that contain pigments. Step 2: These chromatophores allow chameleons to expand or contract, changing the amount of pigment visible on their skin. Step 3: This color-changing ability helps chameleons with camouflage, communication, and temperature regulation. Therefore, the answer is that they change colors for camouflage, communication, and temperature regulation."
Table 14 reveals that Summary3 strongly predicts "cam" (99.61%, likely "camouflage"), while Summary4-6 consistently favor "reg" (likely "regulation") with "communication" as a competing alternative. This demonstrates the Summary Tokens’ ability to extract and compress the three key biological functions (camouflage, communication, temperature regulation) from the detailed mechanistic explanation.
Context-Aware Summarization. Interestingly, Summary2 exhibits context-dependent behavior: predicting "answer" for mathematical and Misconception questions but "Reason" for the scientific explanation. This suggests that Summary2 may encode the question type or intent, asking "What is the nature of this question?" in a single token representation. Such adaptive behavior indicates that the Summary Tokens operate hierarchically, with early tokens capturing meta-information about the response structure and later tokens focusing on specific content.
These empirical findings provide strong evidence that our special tokens function as Summary Tokens, systematically compressing the model’s reasoning conclusions into interpretable representations. The consistent emergence of conclusion-relevant predictions across diverse question types, despite no direct supervision of the prediction layer during training, suggests that the learned hidden representations naturally encode semantic summaries of the model’s intended responses.
Appendix G Analysis of Response Sampling Number Effects
G.1 Experimental Setup
To investigate the relationship between the number of sampled responses and performance across different consistency methods, we conduct experiments with varying response numbers (5, 10, and 20 responses) on four benchmark datasets: GSM8K, MATH, TriviaQA, and TruthfulQA. We compare our proposed Latent Self-Consistency (LSC) method against established baselines including Self-Consistency, WUCS, and USC.
G.2 Results
Table 15 presents the comprehensive results of our analysis across different response sampling numbers. The results reveal several important patterns regarding the scalability and effectiveness of different consistency methods.
G.3 Analysis and Discussion
Effective scaling in factual QA tasks. Table 15 demonstrates that as the number of responses increases, both Self-Consistency (SC) and our proposed Latent Self-Consistency (LSC) show substantial performance improvements in factual QA tasks (GSM8K, MATH, TriviaQA). This indicates that both methods effectively leverage the benefits of increased response numbers that SC originally provides.
Limited scaling in other methods. In contrast, USC and WUCS exhibit minimal performance gains despite the increased number of responses. This limited scaling suggests that these surface-level consistency method and prompt-based method fail to fully capitalize on the fundamental advantage of Self-Consistency—namely, the improved accuracy that comes from having more number of candidate responses to choose from. Our LSC method, however, successfully inherits and utilizes this key benefit of SC.
Consistent performance in open-ended tasks. For the open-ended TruthfulQA task, our LSC method maintains consistently high performance around 94% consistency across different response counts. This stability demonstrates the robustness of our approach in handling more complex, narrative-style responses.
Dynamic TopK limitations in open-ended settings. We observe a notable performance drop in the Dynamic TopK variant when the number of responses reaches 20 for TruthfulQA consistency. This degradation appears to stem from the inherent complexity of open-ended responses. While factual QA tasks produce relatively simple representations that remain distinguishable even with larger response sets, open-ended tasks generate more complex and nuanced representations. This complexity can lead to ambiguous boundaries during majority detection, causing the Dynamic TopK method to make suboptimal boundary decisions.
Figure 17: Accuracy and consistency performance of LLaMA3-8B-Instruct across different summary token counts on the MATH dataset. Accuracy shows no clear trend across token counts, while consistency exhibits a modest upward trend with occasional instability (notably at 8 tokens). Overall, token count has limited impact on performance, allowing selection based primarily on computational efficiency considerations.
Appendix H Impact of Summary Token Count on Model Performance
We investigate the effect of varying summary token counts on model performance to determine the optimal configuration for our approach. This analysis provides insights into the sensitivity of our method to the number of summary tokens used for response representation.
H.1 Performance Analysis
Accuracy Trends: The accuracy metric shows relatively stable performance across different token counts with no clear directional trend. Values fluctuate within a narrow range (0.370-0.376), suggesting that the number of summary tokens does not significantly impact the model’s ability to predict correct final answers.
Consistency Patterns: The consistency metric exhibits a modest upward trend with increasing token count, though with notable fluctuation at 8 tokens before recovering at 10 tokens. While the overall improvement is gradual, this suggests that additional summary tokens may provide slightly enhanced stability in response evaluation, albeit with occasional training instability at higher counts.
Overall Performance Impact: Importantly, our results indicate that summary token count does not dramatically affect model performance. The relatively modest differences across configurations suggest that our approach is reasonably robust to this architectural choice, providing flexibility in selecting token counts based on computational constraints rather than performance requirements alone.
H.2 Computational and Practical Considerations
Given the limited performance sensitivity to token count, computational efficiency becomes a primary consideration in selecting the optimal configuration. Each additional token increases both training and inference overhead without proportional performance gains.
Optimal Configuration Selection: Based on the balance between slight consistency improvements and computational efficiency, we select 6 summary tokens for our main experiments. This configuration provides modest gains over the single-token baseline while avoiding the training instability observed at 8 tokens and maintaining reasonable computational overhead.
Appendix I Summary Token Representation Strategy: Last Token vs Mean Pooling
In-domain Out-of-domain Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MC1 CommonsenseQA LLaMA3.1 8B-Inst. Vanilla (1 path) $83.2\pm 0.2$ $39.4\pm 0.5$ $62.4\pm 0.6$ $68.8\pm 1.4$ $57.2\pm 1.6$ $74.9\pm 3.0$ Self Consistency $\underline{92.2}\pm 0.3$ $52.5\pm 0.2$ $72.7\pm 0.7$ $73.6\pm 0.5$ $\mathbf{63.0}\pm 0.9$ $\mathbf{79.4}\pm 0.8$ WUCS $89.8\pm 0.3$ $45.6\pm 0.8$ $73.2\pm 0.6$ $72.8\pm 1.3$ $61.8\pm 0.6$ $76.6\pm 0.8$ USC $86.2\pm 0.7$ $44.8\pm 1.3$ $70.1\pm 1.1$ $71.1\pm 0.1$ $59.6\pm 0.7$ $74.6\pm 0.8$ LSC (Ours) $\mathbf{92.3}\pm 0.3$ $\underline{52.6}\pm 0.4$ $\underline{74.1}\pm 0.6$ $\underline{73.8}\pm 0.4$ $62.1\pm 0.4$ $79.0\pm 0.5$ +Dynamic TopK $\underline{92.2}\pm 0.2$ $\mathbf{52.9}\pm 0.2$ $\mathbf{75.7}\pm 0.2$ $\mathbf{73.9}\pm 0.4$ $\underline{62.7}\pm 1.1$ $\underline{79.1}\pm 0.8$ Qwen3-8B Vanilla (1 path) $92.8\pm 0.3$ $71.9\pm 0.8$ $48.1\pm 1.2$ $76.7\pm 0.4$ $72.6\pm 0.3$ $81.6\pm 0.7$ Self Consistency $\mathbf{94.5}\pm 0.2$ $\mathbf{76.7}\pm 0.3$ $\mathbf{56.1}\pm 0.5$ $\mathbf{77.2}\pm 0.4$ $\underline{75.2}\pm 0.3$ $\underline{82.6}\pm 0.5$ WUCS $93.8\pm 0.3$ $72.1\pm 0.6$ $49.9\pm 0.7$ $76.8\pm 0.2$ $75.0\pm 0.3$ $81.9\pm 0.6$ USC $93.7\pm 0.2$ $75.6\pm 0.5$ $50.9\pm 1.3$ $76.0\pm 0.5$ $74.4\pm 0.2$ $81.8\pm 0.2$ LSC (Ours) $\mathbf{94.5}\pm 0.1$ $75.9\pm 0.2$ $\underline{55.5}\pm 0.8$ $\underline{77.1}\pm 0.4$ $74.9\pm 0.3$ $\mathbf{82.7}\pm 0.2$ +Dynamic TopK $94.4\pm 0.2$ $\underline{76.6}\pm 0.7$ $53.2\pm 1.0$ $77.0\pm 0.3$ $\mathbf{75.4}\pm 0.3$ $\underline{82.6}\pm 0.4$
Table 16: Standard deviations of performance scores for short-answer tasks across different models and decoding methods. Results show consistently low variance across all approaches on both in-domain and out-of-domain datasets. All values are reported as percentages with standard deviations over three random seeds.
Models Decoding Methods HumanEval HumanEval+ MBPP MBPP+
LLaMA3.1- 8B-Instruct Vanilla (1 path) $54.47\pm 2.2$ $47.13\pm 5.1$ $64.83\pm 2.3$ $54.50\pm 2.0$
WUCS $55.50\pm 2.7$ $51.00\pm 1.2$ $67.03\pm 0.6$ $55.10\pm 0.9$
USC $56.93\pm 2.7$ $50.80\pm 2.9$ $68.43\pm 1.2$ $58.57\pm 1.6$
LSC (Ours) $\underline{57.93\pm 1.9}$ $\underline{51.23}\pm 2.2$ $\underline{69.23}\pm 1.9$ $\underline{58.90}\pm 1.3$
  +Dynamic TopK $\mathbf{58.50}\pm 1.3$ $\mathbf{51.80}\pm 2.1$ $\mathbf{69.30}\pm 0.7$ $\mathbf{60.60}\pm 0.6$
Qwen3-8B Vanilla (1 path) $77.80\pm 2.0$ $73.40\pm 2.2$ $81.03\pm 0.9$ $69.23\pm 3.3$
WUCS $\mathbf{80.30}\pm 3.4$ $\mathbf{75.20}\pm 2.0$ $\underline{82.43}\pm 1.3$ $\underline{70.17}\pm 3.1$
USC $\underline{78.03}\pm 1.4$ $73.17\pm 2.0$ $\mathbf{82.97}\pm 2.6$ $\mathbf{71.00}\pm 1.8$
LSC (Ours) $\underline{79.30}\pm 2.2$ $\underline{74.20}\pm 1.9$ $81.20\pm 2.9$ $69.40\pm 1.2$
  +Dynamic TopK $78.57\pm 2.9$ $73.50\pm 1.6$ $82.00\pm 2.1$ $70.10\pm 2.4$
Table 17: Standard deviations of performance scores for coding tasks. Results show higher variance compared to short-answer tasks, with standard deviations ranging from 0.6 to 5.1 percentage points. All values are reported as percentages with standard deviations over three random seeds.
Figure 18: Accuracy and consistency of the LLaMA3-8B-Instruct model under different aggregation strategies. Mean aggregation consistently outperforms last-token aggregation across both metrics.
A critical design decision in our approach involves how to utilize the multiple summary tokens for final response representation. We investigate two primary strategies: using only the last summary token’s representation versus employing mean pooling across all summary tokens.
I.1 Experimental Setup
We conduct comparative experiments using LLaMA3-8B-Instruct to evaluate the effectiveness of different representation strategies:
  • Last Token Strategy: Utilizes only the final summary token’s hidden representation as the response embedding for contrastive learning and consistency scoring.
  • Mean Pooling Strategy: Computes the average of all summary token representations to create a unified response embedding.
Both approaches are trained under identical conditions to ensure fair comparison, with all other hyperparameters and training procedures remaining constant.
I.2 Performance Comparison
Figure 18 presents the comparative performance of both representation strategies across accuracy and consistency metrics. The results demonstrate a clear advantage for the mean pooling approach over the last token strategy.
Superior Performance of Mean Pooling: The mean pooling strategy consistently outperforms the last token approach across both evaluation metrics. This improvement suggests that aggregating information from multiple summary tokens provides a more robust and comprehensive response representation than relying solely on the final token.
Representational Capacity Analysis: The superior performance of mean pooling can be attributed to its expanded representational space. By incorporating information from multiple tokens, mean pooling effectively increases the dimensionality and richness of the response embedding, allowing the model to capture more nuanced aspects of response content and quality.
I.3 Implementation Decision
Based on these experimental findings, we adopt the mean pooling strategy for all main experiments. This choice maximizes the utilization of our multi-token summary architecture while providing superior performance across key evaluation metrics. The consistent improvement observed across different evaluation scenarios reinforces the robustness of this design decision.
Appendix J Standard Deviations of Main Experiments
As reported in Table 16, standard deviations for short-answer tasks are small (0.1 to 3.0), indicating stable performance across random seeds. In contrast, Table 17 shows that coding tasks—treated as long-form generation—exhibit substantially larger variability. While all multiple-response methods consistently outperform the vanilla baseline, the gaps among those methods are narrow, which prevents a definitive ranking.
Appendix K Difficulty of Short- vs Long-form Question Classification
To demonstrate LSC’s robustness across both short-answer and long-form settings, we conducted an auxiliary experiment that first attempted to classify each incoming question as short-answer or long-form before applying a decoding strategy. The underlying intuition is that if lightweight, efficient methods such as Self-Consistency (SC) or WUCS already perform near-optimally within their respective regimes, one could in principle route a question to the appropriate method after classification. However, this pipeline introduces additional overhead and, as we demonstrate below, the classification itself is nontrivial. Even when the question type is (imperfectly) inferred, LSC remains compelling because it delivers substantially higher consistency than WUCS without requiring prior knowledge of the format.
For classification we employed the commercially strong and efficient model GPT-4o-mini. We curated 50 frequently asked “What”-prefixed questions that often demand elaborated explanations—i.e., they superficially resemble short-answer prompts but are semantically long-form (conceptual definitions, mechanisms, etc.). The full list appears in Figure 20. GPT-4o-mini was prompted to label each question as either short-answer or long-form. It identified only 7 out of 50 as long-form, yielding an effective recognition rate of 14%. By contrast, on more conventional question types (e.g., TriviaQA-style queries), GPT-4o-mini achieves over 90% accuracy in format discrimination. This gap indicates that distinguishing short-answer from long-form questions in the conceptual regime is itself a challenging classification problem. The prompt used for this task is shown in Figure 11.
These findings emphasize two points. First, relying on a separate classification step to select between methods adds complexity and potential failure modes. Second, LSC’s ability to operate without such prior classification while achieving superior consistency compared to WUCS makes it a more streamlined and reliable approach across heterogeneous question formats.
Short- vs. Long-form Question Classification Prompt System: You are an assistant that determines whether a given question requires a short answer or a long answer. User:
Analyze the following question and determ-
ine whether it requires a short answer
form or an open-ended form.
Question: {Your Question Here}
If it requires a short answer, respond
with "#### short". If it requires an open-
ended answer, respond with "#### long".
Figure 19: Prompt used to classify each question as requiring a short-answer or long-form response. The model is instructed to reply with ”#### short” for short-answer format and ”#### long” for open-ended format.
# Question # Question # Question 1 What is Singular Value Decomposition? 2 What is the central limit theorem? 3 What is recursion in computer science? 4 What is quantum entanglement? 5 What is the second law of thermodynamics? 6 What is Bayesian inference? 7 What is a Nash equilibrium? 8 What is CRISPR-Cas9 gene editing? 9 What is a blockchain? 10 What is the Doppler effect? 11 What is a Fourier transform? 12 What is cloud computing? 13 What is entropy in information theory? 14 What is reinforcement learning? 15 What is the Heisenberg uncertainty principle? 16 What is a Monte Carlo simulation? 17 What is an eigenvector? 18 What is the greenhouse effect? 19 What is Big-O notation? 20 What is the placebo effect? 21 What is the Internet of Things? 22 What is public-key cryptography? 23 What is the Turing test? 24 What is plate tectonics? 25 What is convolution in deep learning? 26 What is the twin paradox? 27 What is containerization with Docker? 28 What is a Markov chain? 29 What is genetic drift? 30 What is the Schwarzschild radius? 31 What is agile software development? 32 What is graphene? 33 What is the Pauli exclusion principle? 34 What is a sigmoid activation function? 35 What is SMTP? 36 What is the Monty Hall problem? 37 What is herd immunity? 38 What is the photovoltaic effect? 39 What is backpropagation? 40 What is the Bretton Woods system? 41 What is the Prisoner’s Dilemma? 42 What is a qubit? 43 What is inflation targeting? 44 What is the anthropic principle? 45 What is a RESTful API? 46 What is the difference between HTTP and HTTPS? 47 What is the Fibonacci sequence? 48 What is a deadlock in concurrent programming? 49 What is MRI imaging based on? 50 What is the Pareto distribution?
Figure 20: List of prompt questions used in the study.